OpenStudy (anonymous):

Express y as a function of x. What is the domain? a) log 3 + log y = log (x+2) - log x

OpenStudy (vishweshshrimali5):

log a + log b = log(ab)

OpenStudy (lgbasallote):

i strongly suggest you put everything as one log first

OpenStudy (vishweshshrimali5):

and log a - log b = log (a/b)

OpenStudy (anonymous):

log(y)=log(x+2)-log(x)-log(3) log(y)=log(x+2)-(log(x)+log(3)) log(y)=log(x+2)-(log3x) log(y)=log((x+2)/3x) exponentiating both sides y = (x+2)/3x domain is all real x except x =0 i think im wrong

OpenStudy (vishweshshrimali5):

log (3y) = log (\(\cfrac{x+2}{x}\))

OpenStudy (vishweshshrimali5):

3y = \(\cfrac{x+2}{x}\) y = \(\cfrac{x+2}{3x}\)

OpenStudy (anonymous):

Extremity you're close the answer is: {x< -2 or x > 0}

OpenStudy (anonymous):

I'm confused about the domain though..

OpenStudy (anonymous):

oh whoops, dangerous (x+2)/3x > 0 , since log has to always be positive that means (x+2)3x>0 (x+2)x>0 graph a parabola and you see that x<-2 and x>0 point is, when you make y the subject, you remove a fundamental analysis on the equation. that is, the presence of the log, and so the domain is what it is